Hydra Report - March 2004: Problems encountered Hydra block

Updated on June 15, 2021, 7:49 am

  1. CCD dewar seems to require many refills before reaching correct temperature. Vacuum indicator appears normal, but perhaps we don't have a fine enough instrument to indicate that there is outgassing.

    Ramon will check out the vacuum and prepare a new molecular sieve during the week of 21/3
     

  2. Observers reported problems with the rotation for their fields near the pole (see email from Peter Frinchaboy and Bill Kunkel). The Hydra rotation was measured on the engineering night at the equator, but x_offset and y_offset were left at 0.

    On the next engineering night, we will measure the rotation near the equator and adjust x_offset and y_offset by going to the pole.

    Rolando is looking at exactly how big the axial alignment error is.

    Can Roberto check the optical alignment soon?
     

  3. Minor software problem: Locked FOPS 120 always activated when turn on guider.

    Rolando will add a fix to the code

  4. Lots of trouble setting up echelle, caused basically by very weak and hard to identify ThAr lines and the fact that observers were operating in UV

    First, Hernan, Ricardo, Knut, Daniel, and Patricio will write up a detailed set of procedures, hints, and tricks for doing the echelle setup

    We want to investigate getting replacement ThAr lamps--Oscar will contact Kitt Peak to find out what they are using

     

  5. The Ar penray is actually Hg, the Ne penray appears weak

    Observer support will check out what is currently in the chimney.

    Oscar and Knut will contact Kitt Peak about what lamps they have and who supplies them

    Hernan, Oscar, Ricardo, Knut, Patricio will also make a web page listing all of our comparison lamp supplies

     

  6. Some gripper problems: short circuit in z-limit switch, and one instance where the status of the limit switches (seen through swchart and inputchart commands) were changing erratically

    Andres and Ricardo will check out the z-limit switch and the general mechanical tuning of the gripper during the week of 21/3. Andres will also see whether a new z-limit switch design may be implemented before April 16, 2004.

    The trouble with the erratic limit switch signals was solved by replacing the spare Galil with the original Galil. Trouble is that the original Galil had some problems of its own, such as getting lost and producing "x-y stage not at destination" errors. The Galil boxes have been sent back to the company twice for repair, but have been returned with the statement that there is nothing wrong with them. Gale will ask Galil for a quote on a new replacement.
     

  7. At the start of the Hydra block, the guider parameters were set incorrectly

    FIXED after David changed the values and Rolando incorporated the fix into automated software.